Write a function rule for the table of values; x=0,1,2,-3 y=1,3,5,-5
Answer by venugopalramana(3286)

SINCE THERE IS NO HINT ON THE NATURE OF RELATION,WE HAVE TO GO BY TRIAL AND ERROR
LET US TRY LINEAR RELATION..THAT IS Y=AX+B
AT X=0,Y=1....SO A*0+B=1...OR B=1
SO WE HAVE Y=AX+1
AT X=1,Y=3....SO A*1+1=3...ORA=3-1=2
SO WE HAVE Y=2X+1
NOW LET US TET FOR OTHER CASES WHETHER IT HOLDS
AT X=2,Y=5....2*2+1=5.....OK
AT X=-3,Y=-5.......2*-3+1=-6+1=-5.....OK
HENCE THE RELATION IS Y=2X+1